Circular No. 7634

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 K4 (SOHO) Further to IAUCs 7613 and 7631, D. A. Biesecker has reported mea-

surements by D. Hammer of another comet found by Biesecker on LASCO C2 images obtained with the SOHO spacecraft. The reduced observations and orbital elements by B. G. Marsden appear on MPEC 2001-K27.

2001 UT  α_{2000}  δ_{2000}
May 22.436  3^{h}58.^{m}3  +18^{o}47′

COMET C/2001 K5 (LINEAR) An object reported as asteroidal by the LINEAR program on May 17

(with prediscovery LINEAR observations on Apr. 30 identified by G. V. Williams) and posted on The NEO Confirmation Page has been found to be slightly diffuse with coma diameter 8′′ on CCD images taken at Kleť on May 27.0 UT by M. Tichý and J. Tichá and to be strongly condensed with a 12′′ coma and a 13′′ tail in p.a. 210^{o} on 300-s R-band exposures taken with the 1.8-m f/1 Vatican Advanced Technology Telescope at Mt. Graham on May 27.3 by C. W. Hergenrother, T. B. Spahr, and M. Nelson.

2001 UT  α_{2000}  δ_{2000}  m_{2}
May 17.27721  15^{h}28^{m}10.^{s}02  −24^{o}47′39′′.1  16.7

All available astrometry and the following parabolic orbital elements appear on MPEC 2001-K40.

T = 2002 Oct. 11.586 TT  ω = 46.^{o}950
Ω = 237.461 }2000.0
q = 5.19269 AU  i = 72.691
2001TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 21 15^{h}26.^{m}10 −24^{o}20.^{′}3 5.408 6.413 172.^{o}8 1.^{o}1 14.2
26 15 23.36 −23 43.1 5.396 6.392 168.6 1.8 14.2
31 15 20.70 −23 05.4 5.393 6.372 163.5 2.6 14.2
June 5 15 18.16 −22 27.5 5.397 6.351 158.3 3.4 14.2
10 15 15.78 −21 49.8 5.409 6.330 152.9 4.2 14.2
15 15 13.57 −21 12.4 5.429 6.310 147.6 4.9 14.2
20 15 11.57 −20 35.7 5.455 6.290 142.3 5.7 14.2
25 15 09.79 −19 59.8 5.488 6.270 137.0 6.4 14.2
30 15 08.26 −19 25.1 5.527 6.249 131.7 7.0 14.2
2001 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green